# DeID-GPT: Zero-shot Medical Text De-Identification by GPT-4

### v2

@@ III. Datasets
The i2b2/UTHealth Challenge: We benchmark our proposed method using the 2014 i2b2/UTHealth de-identification challenge dataset  88. Upon request, the Blavatnik Institute of Biomedical Informatics at Harvard University granted us access to this dataset. This dataset contains 1,304 free-form clinical notes of 296 diabetic patients. All PHI entities were manually annotated and replaced with surrogates. Specifically, names, professions, locations, ages, dates, contacts and IDs were replaced by surrogate information to protect privacy and facilitate de-identification research. For example, if there is a real patient named ”Mr. James McCarthy” who visited the hospital on 12/01/2013, these strings will be replaced by ”Mr. Joshua Howard” and ”04/01/2060”, respectively. Figure 4 presents a sample clip of this data, and figure 5 presents the same excerpt de-identified by ChatGPT/GPT-4.
@@ A Simple and Versatile Framework
Specifically, in-context learning eliminates the need for fine-tuning, thereby saving significant amounts of time and reducing pipeline complexity  89; 90. Supervised fine-tuning is not a trivial task and requires sufficient experience in NLP.
In addition, this approach has exceptional zero-shot and few-shot learning capabilities  45; 90; 91, which can reduce the need for large-scale annotation efforts, as the model can learn from a relatively small amount of annotated data.
By incorporating contextual information, which allows the model to better understand the meaning and context of the data it is processing, in-context learning enables the adaptation to new tasks and domains with great ease  45; 90. Indeed, LLMs are very versatile and flexible. Unlike any previous methods, no code or procedural changes are required when applying our solution to different hospitals, different languages and different data formats.
@@
Initialize: Initialized ChatGPT $model$
@@
In this template, we first specify the task for the LLM to complete. It is necessary to provide the task information upfront, since it is easy for the user to comprehend and explain the task to be executed by the model. In addition, given the autoregressive nature (i.e., the next token is generated based on previous tokens) of the GPT-based models  46, it is important to push this information upfront for the LLM to better understand the user’s demands and generate the most relevant results.
@@
Prompt design is an engineering process that combines art and science  92; 93. Since this is the first study on utilizing LLMs to perform de-identification and the first work in exploring the ability to de-identify clinical data, we believe it is necessary to share our experience with both good examples and counterexamples. We have presented the optimal prompt template in the previous section. In this section, we present ineffective prompts that lead to sub-optimal results. The list of ineffective prompts is non-exhaustive, and these samples are exemplary caveats.
@@ V. Experiment & Results
We compare ChatGPT and GPT-4 against several baselines methods, BERT  43, RoBERTa  44, ClinicalBERT  94.
BERT is the classic transformer-based language model  43. The bi-directional transformer architecture enables impressive contextual understanding, and BERT has quickly become one of the most popular NLP models. It has been widely used in many domains and industries, and there are various offshoot models that are equipped with improved architectural design  95 or are trained on domain-specific data  96.
RoBERTa improves from BERT by removing the next sentence prediction (NSP) objective  44. It is also trained with larger batches and on more training data. In addition, the masking patterns in RoBERTa are dynamically changed. Overall, RoBERTa can be regarded as a more refined version of the BERT model.
ClinicalBERT is a BERT-based model further pre-trained on the Medical Information Mart for Intensive Care III (MIMIC-III) dataset  97, a large collection of 2,083,180 clinical notes from the Beth Israel Deaconess Medical Center. This database contains data of 38,597 patients (admitted between 2001 and 2012) and is commonly used in clinical NLP  98. The ClinicalBERT model is therefore a suitable model for medical text processing.
@@ VI. Discussion
Despite the impressive performance of ChatGPT and GPT-4 for data de-identification, these models could only be accessed through online APIs, making them impossible to be applied in a hospital setting as patient data cannot be stored nor transmitted to a non-authorized external party. Furthermore, as the code and implementation details of ChatGPT/GPT-4 are not open to the public, we cannot fully validate their functionalities to ensure these models are HIPAA-compliant and meet the standards of the hospital’s quality management system (QMS). Therefore, it would be vital to use open-source or in-house trained LLMs for the de-identification task and deploy them locally to ensure data security, privacy, integrity, and proper adherence to HIPAA guidelines. Thus, we are investigating solutions for the locally-deployed DeID-GPT. For example, we experimented with LLaMA  99 in this work to preliminarily examine its performance. However, LLaMA failed to generate coherent and relevant outputs for all test cases, and is therefore not directly comparable to other tested methods. We also intend to try other open-source LLMs such as OPT  48 and BLOOM  100. These models are open-sourced and were trained with data on the public domain, making it possible to reuse without licensing concerns  99. To deal with the challenge of limited computational resources at the local site, we will also explore the state-of-the-art model quantization  101, acceleration  102 and GPU parallelization  103; 104 techniques to streamline the deployment.
@@
To the best of our knowledge, there are no LLMs specifically tailored to the medicine and healthcare domain with medical text such as the clinical notes and radiology reports used in this work, let alone for more diversified medical specialties. Given the success of previous domain-specific language models such as BioBERT  105 and BioGPT  106, we believe it is promising and helpful to develop domain-specific LLMs. At the data level, it is possible to continue the pre-training process on domain-specific data. This is a proven route to success since it is well known that language models can perform better on domain-specific benchmarks and tasks when trained with relevant data and knowledge  96; 107 as the models are exposed to domain vocabulary, jargons, terms and writing styles. At the architecture level, it is also possible to modify and optimize LLMs for specific domains. For example, it is possible to design modules that can more effectively capture sensitive HIPAA-protected information by prioritizing such information in the input  108. It might also be insightful to explore the possibility of efficient passing of inputs with sparsely activated forward pass blocks  109 to achieve performance or efficiency gains.
@@ VII. Conclusion and Future Perspective
As far as we know, this study is the first work in the NLP literature to investigate the possibility of employing LLMs such as ChatGPT and GPT-4 for data de-identification. In particular, we partnered with clinicians to evaluate these powerfully versatile modern models on medical text anonymization. Experimental results indicate that ChatGPT and GPT-4 have excellent abilities in de-identifying medical data compared to other LLMs. The application of LLMs to medical text data has already shown promise in providing valuable insights into various medical conditions and diseases, and there is significant potential for future integration with other modalities such as medical imaging data 110.

### v3

@@ III. Datasets
The i2b2/UTHealth Challenge: We benchmark our proposed method using the 2014 i2b2/UTHealth de-identification challenge dataset  stubbs2015annotating. Upon request, the Blavatnik Institute of Biomedical Informatics at Harvard University granted us access to this dataset. This dataset contains 1,304 free-form clinical notes of 296 diabetic patients. All PHI entities were manually annotated and replaced with surrogates. Specifically, names, professions, locations, ages, dates, contacts and IDs were replaced by surrogate information to protect privacy and facilitate de-identification research. For example, if there is a real patient named ”Mr. James McCarthy” who visited the hospital on 12/01/2013, these strings will be replaced by ”Mr. Joshua Howard” and ”04/01/2060”, respectively. Figure 4 presents a sample clip of this data, and figure 5 presents the same excerpt de-identified by ChatGPT/GPT-4.
@@ A Simple and Versatile Framework
Specifically, in-context learning eliminates the need for fine-tuning, thereby saving significant amounts of time and reducing pipeline complexity  qin2023chatgpt; zhou2023comprehensive. Supervised fine-tuning is not a trivial task and requires sufficient experience in NLP.
In addition, this approach has exceptional zero-shot and few-shot learning capabilities  brown2020language; zhou2023comprehensive; bang2023multitask, which can reduce the need for large-scale annotation efforts, as the model can learn from a relatively small amount of annotated data.
By incorporating contextual information, which allows the model to better understand the meaning and context of the data it is processing, in-context learning enables the adaptation to new tasks and domains with great ease  brown2020language; zhou2023comprehensive. Indeed, LLMs are very versatile and flexible. Unlike any previous methods, no code or procedural changes are required when applying our solution to different hospitals, different languages and different data formats.
@@
Initialize: Initialized ChatGPT modelmodel
@@
In this template, we first specify the task for the LLM to complete. It is necessary to provide the task information upfront, since it is easy for the user to comprehend and explain the task to be executed by the model. In addition, given the autoregressive nature (i.e., the next token is generated based on previous tokens) of the GPT-based models  radford2018improving, it is important to push this information upfront for the LLM to better understand the user’s demands and generate the most relevant results.
@@
Prompt design is an engineering process that combines art and science  reynolds2021prompt; jiang2022promptmaker. Since this is the first study on utilizing LLMs to perform de-identification and the first work in exploring the ability to de-identify clinical data, we believe it is necessary to share our experience with both good examples and counterexamples. We have presented the optimal prompt template in the previous section. In this section, we present ineffective prompts that lead to sub-optimal results. The list of ineffective prompts is non-exhaustive, and these samples are exemplary caveats.
@@ V. Experiment & Results
We compare ChatGPT and GPT-4 against several baselines methods, BERT  devlin2018bert, RoBERTa  liu2019roberta, ClinicalBERT  alsentzer2019publicly.
BERT is the classic transformer-based language model  devlin2018bert. The bi-directional transformer architecture enables impressive contextual understanding, and BERT has quickly become one of the most popular NLP models. It has been widely used in many domains and industries, and there are various offshoot models that are equipped with improved architectural design  rogers2021primer or are trained on domain-specific data  gu2021domain.
RoBERTa improves from BERT by removing the next sentence prediction (NSP) objective  liu2019roberta. It is also trained with larger batches and on more training data. In addition, the masking patterns in RoBERTa are dynamically changed. Overall, RoBERTa can be regarded as a more refined version of the BERT model.
ClinicalBERT is a BERT-based model further pre-trained on the Medical Information Mart for Intensive Care III (MIMIC-III) dataset  johnson2016mimic, a large collection of 2,083,180 clinical notes from the Beth Israel Deaconess Medical Center. This database contains data of 38,597 patients (admitted between 2001 and 2012) and is commonly used in clinical NLP  pandey2022comprehensive. The ClinicalBERT model is therefore a suitable model for medical text processing.
@@ VI. Discussion
Despite the impressive performance of ChatGPT and GPT-4 for data de-identification, these models could only be accessed through online APIs, making them impossible to be applied in a hospital setting as patient data cannot be stored nor transmitted to a non-authorized external party. Furthermore, as the code and implementation details of ChatGPT/GPT-4 are not open to the public, we cannot fully validate their functionalities to ensure these models are HIPAA-compliant and meet the standards of the hospital’s quality management system (QMS). Therefore, it would be vital to use open-source or in-house trained LLMs for the de-identification task and deploy them locally to ensure data security, privacy, integrity, and proper adherence to HIPAA guidelines. Thus, we are investigating solutions for the locally-deployed DeID-GPT. For example, we experimented with LLaMA  touvron2023llama in this work to preliminarily examine its performance. However, LLaMA failed to generate coherent and relevant outputs for all test cases, and is therefore not directly comparable to other tested methods. We also intend to try other open-source LLMs such as OPT  zhang2022opt and BLOOM  scao2022bloom. These models are open-sourced and were trained with data on the public domain, making it possible to reuse without licensing concerns  touvron2023llama. To deal with the challenge of limited computational resources at the local site, we will also explore the state-of-the-art model quantization  dettmers2022llm, acceleration  zhang2020accelerating and GPU parallelization  narayanan2021efficient; zeng2022acctfm techniques to streamline the deployment.
@@
To the best of our knowledge, there are no LLMs specifically tailored to the medicine and healthcare domain with medical text such as the clinical notes and radiology reports used in this work, let alone for more diversified medical specialties. Given the success of previous domain-specific language models such as BioBERT  lee2020biobert and BioGPT  luo2022biogpt, we believe it is promising and helpful to develop domain-specific LLMs. At the data level, it is possible to continue the pre-training process on domain-specific data. This is a proven route to success since it is well known that language models can perform better on domain-specific benchmarks and tasks when trained with relevant data and knowledge  gu2021domain; rezayi2022agribert as the models are exposed to domain vocabulary, jargons, terms and writing styles. At the architecture level, it is also possible to modify and optimize LLMs for specific domains. For example, it is possible to design modules that can more effectively capture sensitive HIPAA-protected information by prioritizing such information in the input  liao2023mask. It might also be insightful to explore the possibility of efficient passing of inputs with sparsely activated forward pass blocks  fedus2021switch to achieve performance or efficiency gains.
@@ VII. Conclusion and Future Perspective
As far as we know, this study is the first work in the NLP literature to investigate the possibility of employing LLMs such as ChatGPT and GPT-4 for data de-identification. In particular, we partnered with clinicians to evaluate these powerfully versatile modern models on medical text anonymization. Experimental results indicate that ChatGPT and GPT-4 have excellent abilities in de-identifying medical data compared to other LLMs. The application of LLMs to medical text data has already shown promise in providing valuable insights into various medical conditions and diseases, and there is significant potential for future integration with other modalities such as medical imaging data wang2023chatcad.